• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name A. Deur

    Log entry time 00:03:54 on November17,2009

    Entry number 300453

    keyword=Swing shift summary

    Shift started with no beam.

    16:10 MCC called saying they will send test and the CW beams in the Hall. Target is in empty position.
    16:50 CW beam in the hall (~90A), but only spikes (MCC still tuning for Compton).
    17:20 MCC called. Tune beam will be sent to the Hall.
    19:15. Alexandre C. called saying he was mostly done with the Compton tune and we can start beam recovery procedure. Check that beam position is good. We requested a Harp Scan.
    19:40 MCC called saying that they can't get any of the Hall A harps to move.
    19:52 After talking with RC, we decided to bypass for now the harp scan and go ahead with the next steps of beam recovery. Meanwhile, they are paging an expert to look into the harp problem.
    20:32 Beam looks good. All steps completed except the Harp.
    20:34 Requested a Control Access for Alexandre C. and possibly to reboot the Harp control.
    20:50 Access done. We are requesting 2 uA to test Compton run and do the fast beam recovery
    21:15. Beam back, 2 uA, VDC HV off.
    21:25. Taking data during the Compton rate check. Production but at low current. Parity run 14002, LHRS 25542, RHRS 5090.
    21:33 beam down to reboot an IOC (may fix the harp problem)
    21:50 Harps seem to be fixed. We move to empty target position for Harp scan.
    22:05 Harp scan done. Beam profile looks good: 190*160 on HO3a and 190*150 on HO3b.
    Continuing runs on empty target for Compton rate checking.
    23:18 Compton test done. Request controlled access to put back the shielding.
    23:40 Access done. Request beam at 2 uA for quick recovery procedure check.
    23:53. 2 uA Beam looks good. Request 50 uA for first step production.
    23:56 Everything looks fine. Asked for 100uA.
    Production runs: happex run number 14005 LHRS: 25546 RHRS: 5094